[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jul 13, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Wed Jul 13 15:32:05 UTC 2016


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
-The TAFB shift log scripts are intermittently reporting issues to the TSB  
spreadsheet. TSB will investigate.
HSU:
-TAFB Dvorak fixes came in faster overnight. TSB will continue to monitor  
this issue in case additional tweaks are needed again.
-Himawari data is still out. TSB will investigate.
-TSB is still investigating duplicate shift log emails.
-TSB is still investigating the Surface Wind Field graphics issues.
Other items:
-The Storm Surge AWIPS Sprint continues all week.
-After the HWRF/GFDL upgrades yesterday, the grib2 files are arriving later  
than usual. Note that the new files are much larger than the old files,  
which may account for the latency. TSB is working with NCO to resolve this  
issue.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160713/6b6e99dd/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list